Methylamine (CH_3NH_2) is a weak base. Write the r
Answer #1

a) CH3NH2 + H2O ---> CH3NH3+ + OH-

Kb = [CH3NH3+] [OH-]/[CH3NH2]

b) pH = 11.7 , pOH = 14-11.7 = 2.3

[OH-]= 10^(-2.3) = 0.005 M

0.005 = C*X

0.005 = 0.065*X

X = 0.077

Kb = cx^2

    = 0.065*(0.077)^2


   kb = 3.85*10^(-4)

c) ka = kw/kb = 10^(-14)/(3.85*10^(-4))

    = 2.6*10^-11